Which type of triangle can be constructed with a 50°
angle between two 8-inch sides?

Answer by ewatrrr(24785)


Hi
Which type of triangle can be constructed with a 50°angle between two 8-inch sides?
Isosceles Triangle with base angles of (180-50)/2 = 130/2 = 65°